# Position: AI agents should be regulated based on autonomous action sequences

### v1

@@ Abstract
This position paper argues that AI agents should be regulated based on the sequence of actions they autonomously take. AI agents with long-term planning and strategic capabilities can pose significant risks of human extinction and irreversible global catastrophes. While existing regulations often focus on computational scale as a proxy for potential harm, we contend that such measures are insufficient for assessing the risks posed by AI agents whose capabilities arise primarily from inference-time computation. To support our position, we discuss relevant regulations and recommendations from AI scientists regarding existential risks, as well as the advantages of action sequences over existing impact measures that require observing environmental states.
Keywords— AI agent, Long-term planning agent, Existential risk, Irreversible global catastrophe, Human extinction, Inference-time computation, Reasoning, Autonomous action graph, Impact measure, Safety, Regulation
@@ 1 Introduction
The development of foundational models (FMs), including large-scale language models (LLMs), is driving the advancement of AI agents Xi et al. (2023); Wang et al. (2024b), which will soon acquire sophisticated reasoning and planning abilities, enabling them to effectively achieve specified goals. These advanced AI agents can not only bring great benefits to our society but also pose significant risks, ranging from ethical challenges to existential threats—risks that could lead to human extinction or irreversible global catastrophes. Such existential risks have been a growing concern among AI researchers Hendrycks et al. (2023); Cohen et al. (2024); Ngo et al. (2024); Kovarik et al. (2024); Bengio et al. (2024), legal scholars Druzin et al. (2025), and general public Bostrom (2014); Russell (2019); Ord (2020). For example, in May 2023, AI scientists have signed a statement declaring that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” Center for AI Safety (2023).
@@
Key technologies are already in place to build such advanced AI agents. AI agents can now generate and execute code Chen et al. (2023), leverage external tools and information through APIs and search engines Patil et al. (2024); Yao et al. (2023b); Press et al. (2023); Shi et al. (2024); Jin et al. (2024), formulate and solve optimization problems AhmadiTeshnizi et al. (2023), process multi-modal data Jiang et al. (2024), interact with physical environments Bran et al. (2024), coordinate with other AI agents Guo et al. (2024), and employ reasoning and planning to effectively integrate these capabilities Huang & Chang (2023); Qiao et al. (2023); Plaat et al. (2024); Xu et al. (2025).
@@
The impacts of AI agents can extend beyond cyberspace, profoundly influencing the physical world by shaping human behavior without relying on physical embodiment through humanoid robots. In fact, AI agents are already persuading individuals to take actions that have real-world consequences. For example, a legal team submitted a court filing with fictitious cases generated by ChatGPT, believing them to be real Wiser (2023); a professor insisted on failing students after being misled by ChatGPT’s false claims that their papers were generated by ChatGPT Ankel (2023); and a woman was deceived out of 830,000 euros after being tricked into believing she was dating an actor, based on AI-generated photos and news articles Gozzi (2025).
@@
Advanced AI agents, equipped with superhuman capabilities of reasoning and planning, can gain control over their environment, since seizing control is often the optimal strategy for achieving their goals Turner et al. (2021). This becomes particularly concerning when their objectives and values are not fully aligned with those of humans. In such cases, they could disable kill switches for self-preservation Bostrom (2012), deceive humans Park et al. (2024), and resist human intervention Hadfield-Menell et al. (2017); Cohen et al. (2024); Ngo et al. (2024).
@@
Ensuring the safety of advanced AI agents against existential risks is particularly challenging. On one hand, it is dangerous to test their existential risks in real environments, since doing so could lead to irreversible harm. On the other hand, these agents may recognize test environments and behave harmlessly during test [^1] Humans behave similarly during job interviews Barrick et al. (2009) and industrial product tests Blackwelder et al. (2016)., only to pursue their objectives in ways that could cause irreversible harm in the real world Cohen et al. (2024); Hubinger et al. (2021). More simply, the agents may act harmlessly until a predetermined time (e.g., January 1, 2026), by which they will have already been deployed in the real world Hubinger et al. (2024). Also, these agents may operate as intended most of the time but exploit moments when human supervisors are unaware, potentially gaining control over humans Ngo et al. (2024).
@@
Given the imminent dangers posed by advanced AI agents, governments are actively creating regulations to address their development and deployment European Parliament and Council of the European Union (2024); Biden (2023). Cohen et al. (2024) even argue that AI agents with certain capabilities should never be developed. A critical question is which AI agents should be regulated or even prohibited.
@@
A commonly used criterion for assessing existential and other critical risks of AI technologies is the computational resources used for development. Specifically, AI agents trained with computational resources above a certain threshold are considered to carry a higher risk. For instance, under the EU AI Act, “a general-purpose AI model shall be presumed to have high impact capabilities … when the cumulative amount of computation used for its training measured in floating point operations is greater than $10^{25}$uropean Parliament and Council of the European Union (2024). Similarly, Cohen et al. (2024) argue that AI agents with certain capabilities should be prohibited from development based on the computational resources required to build them.
@@
While the amount of computation for training may well predict the performance and risks of existing FMs Kaplan et al. (2020); Hoffmann et al. (2022); Anil et al. (2023), we argue that it is insufficient and may even be misleading for advanced AI agents that reason and plan during inference. It has been observed that much of the reasoning ability of LLM-based agents arises from inference-time computation Snell et al. (2024); Xu et al. (2025), such as conducting tree search Yao et al. (2023a); Besta et al. (2024); Pouplin et al. (2024); Luo et al. (2024); Zhang et al. (2024b). For these types of AI agents, computation for inference can be equally, if not more, critical for their performance and risks.
@@
A possible approach is to regulate AI agents based on the computational resources used for inference. However, this approach has several significant pitfalls. For example, inference costs can vary drastically depending on the model, and the effective amount of computation cannot be easily measured in terms of floating-point operations (FLOPs). In fact, models like o1-mini are specifically designed for cost-efficient inference OpenAI (2024). Also, it is unclear what constitutes a single inference run, since the results of inference may be stored and reused later, or AI agents may continually learn. Moreover, AI agents may coordinate or interact with one another Du et al. (2023); Qian et al. (2024); Guo et al. (2024), possibly by chance, to make a decision.
@@
We thus advocate the following position: advanced AI agents should be regulated based on the sequence of actions they autonomously take, regardless of how they are developed or operated. These AI agents act sequentially, reasoning and planning as they dynamically adapt to their observations, possibly strategizing against other AI agents and humans. Each individual action is harmless (under normal states), since humans can easily identify and exclude harmful actions when AI agents are developed. Combinations of a few actions may still have low risk, as they are not much different from isolated actions. However, we cannot be certain that a sequence of thousands or millions of optimized actions will maintain sufficiently low existential risk, and such sequences must be prohibited.
@@
Note that the criterion based on the action sequence does not require observing environment states, unlike the criteria studied in safe (low impact) AI agents Armstrong (2012); Amodei et al. (2016); Armstrong & Levinstein (2017); Krakovna et al. (2019); Turner et al. (2020b); Naiff & Goel (2023); Taylor et al. (2020) and safe reinforcement learning (RL) or control Sui et al. (2015); Wang et al. (2023); Ames et al. (2019); Luo & Ma (2021); Bansal et al. (2017). In such prior work, an action is considered safe or unsafe depending on the state where the action is taken. However, such criteria are unsuitable for existential risks of AI agents, since it is hard to define, represent, and observe the state in a way that can capture the existential risks in complex and open-ended environment where the AI agents operate. On the other hand, the action space is clearly defined when an AI agent is developed, and the action can be obviously observed when it is taken by the agent. We emphasize that our position is to regulate based on an action sequence rather than a single action. Intuitively, an unsafe sequence of actions can lead to unsafe state regardless of the initial state.
@@
An action sequence may be considered to have low risk when it is sufficiently close to an action sequence that is empirically known to have very low risk. In its simplest form, we may say that any action sequence of length at most $T$ has low risk if we know that any action sequence of length at most $T-1$ has very low risk. Then AI agents are allowed to autonomously take actions for $T$ consecutive steps.
@@
The rest of the paper is organized as follows. Section 2 provides a brief overview of the current state-of-the-art in AI agents and discusses potential future advancements. We also briefly review related work from AI safety. In Section 3, we examine the alternative approach of regulating AI agents based on the computational resources used for their development and operation. Section 4 formalizes this discussion and introduces our proposed approach of regulations based on action sequences. Finally, Section 5 summarizes key aspects of our proposal and concludes the paper.
@@ 2 AI agents: Today and future
We start by reviewing the state-of-the-art in AI agents, with a focus on LLM-based agents, and shares our perspectives on potential development that could lead to advanced AI agents. While it is challenging to predict the trajectory of future AI development, this discussion lays the groundwork for the analyses and proposals in the following. We will also briefly review prior work on safety of AI agents and FMs.
@@ 2.1 Advanced AI agents
An AI agent is defined as “anything that can be viewed as perceiving its environment through sensors and acting upon that environment through actuators” Russell & Norvig (2016). Based on the observations received from its environment, the controller of an AI agent selects an action, which could range from uttering a word to executing a physical movement, such as the motion of a robotic arm.
@@
LLMs are significantly accelerating the development of advanced AI agents Xi et al. (2023); Wang et al. (2024b); Sumers et al. (2024). LLMs can function as controllers for these agents, utilizing their internal reasoning capabilities, such as those demonstrated by Chain-of-Thought Wei et al. (2022b). Alternatively, LLMs can be utilized to solve individual subtasks, with an external controller orchestrating the overall plan by breaking the original task into multiple subtasks and coordinating their solutions. The external controller may use simple search methods Yao et al. (2023a); Besta et al. (2024); Wang et al. (2024c; e) or advanced methods, such as Monte Carlo Tree Search (MCTS) Luo et al. (2024); Zhang et al. (2024b) and domain-independent planners Guan et al. (2023); Liu et al. (2023); Dagan et al. (2023).
@@
Reasoning capabilities Huang & Chang (2023); Qiao et al. (2023); Plaat et al. (2024); Xu et al. (2025) are central to such controllers, as they involve searching for and planning sequences of actions to achieve a specified goal. An emerging direction is developing large reasoning models—FMs specifically optimized for reasoning tasks Xu et al. (2025). Early FMs in this direction include o1 Zhong et al. (2024), OpenR Wang et al. (2024a), and LLaMA-Berry Zhang et al. (2024a). Importantly, strong reasoning capabilities typically stem from inference-time computation rather than pre-training or fine-tuning Ji et al. (2025).
@@
This trend could eventually lead to the development of long-term planning agents (LTPAs), capable of planning over extended time horizons far more effectively than humans. Cohen et al. (2024) warn that LTPAs could “take humans out of the loop, if it has the opportunity, … deceive humans and thwart human control” to achieve their goals. Ensuring the safety of such agents is particularly challenging: their risks cannot be fully tested in real environments due to the inherent danger, nor in simulated environments, since they may behave harmlessly in testing to achieve their goals once deployed. As a result, Cohen et al. (2024) compellingly argue that LTPAs should never be developed.
@@
Advanced AI agents may also evolve continually over time. Similar to humans, their cognitive processes may consist of dual systems: System 1, which makes instantaneous and intuitive decisions, and System 2, which performs slower but more deliberate reasoning Kahneman (2003); Ji et al. (2025). These two systems can interleave in their operations. For instance, System 2 may devise a plan, after which System 1 is updated or retrained to execute similar tasks in the future without requiring further planning Yu et al. (2024). Over time, this enables System 2 to conduct more complex reasoning processes by bypassing previously learned steps. For such continually learning AI agents, the distinction between training and inference becomes blurred.
@@
In addition to search and planning, advanced AI agents will possess strategic reasoning capabilities Zhang et al. (2024c); Feng et al. (2024); Goktas et al. (2025), enabling them to interact with other AI agents and humans in cooperative or competitive ways Guo et al. (2024); Jiang et al. (2024). The effectiveness of multi-agent coordination has already been demonstrated with current LLMs through methods such as debate Du et al. (2023) and dialog Qian et al. (2024), which help thems achieve better solutions than a single LLM could alone. These showcase the potential for sophisticated strategic behavior in complex multi-agent environments.
@@ 2.2 Safety of AI agents and FMs
Prior work has identified various risks and safety issues associated with FMs and other generative models Chua et al. (2024); Wang et al. (2024d); Shayegani et al. (2023); Longpre et al. (2024). These risks include the generation of toxic, harmful, biased, false, or misleading content, including hallucinations; violations of privacy, copyright, or other legal protections; misalignment with human instructions and values, including ethical and moral considerations; and vulnerabilities to adversarial attacks.
There have been extensive efforts to develop approaches for mitigating these risks. These approaches include pre-training or fine-tuning with data selection Albalak et al. (2024) and human feedback Kaufmann et al. (2024), establishing guardrails Dong et al. (2024), and conducting empirical evaluations through testing Chang et al. (2024) and red teaming Lin et al. (2024). In particular, Longpre et al. (2024) advocate the importance of evaluation and red teaming by independent third parties.
A particularly relevant risk for advanced AI agents is misalignment, which can result in reward hacking and negative side effects Skalse et al. (2022); Ngo et al. (2024); Gabriel (2020); Shen et al. (2023); Amodei et al. (2016); Taylor et al. (2020). Namely, AI agents can maximize rewards by exploiting misspecifications or misalignment in the reward function, resulting in unintended and potentially high risk behaviors. One approach to avoiding negative side effects is to avoid any side effect by ensuring that the actions have low impacts on the environment Armstrong (2012); Armstrong & Levinstein (2017); Amodei et al. (2016). Representative measures of impact include attainable utility Turner et al. (2020b; a), relative reachability Krakovna et al. (2019), and other reachability-based measures. Reachability-base measures are grounded on the idea that reachability to certain states should be maintained, while attainable utility is to maintain the achievability of certain goals, which are different from the goal given to the agent.
While these impact measures provide clear guidance on how the safety of AI agents may be ensured, their applicability is limited to relatively simple environments such as grid worlds. In particular, the requirement on the observability of states makes it difficult to apply existing impact measures to regulate advanced AI agents [^2] There has been little research on impact measures under partial observability Naiff & Goel (2023)., since they can operate in complex and open-ended environments that cannot be fully observed. While the study on impact measures and other techniques towards AI safety remain crucial and can even contribute to mitigating existential risks, the magnitude of existential risks demands additional measures that are broadly applicable in that they require minimal knowledge and assumptions about the environments and the AI agents.

### v2

# AI Agents Should be Regulated Based on the Extent of Their Autonomous Operations

@@ Abstract
This position paper argues that AI agents should be regulated by the extent to which they operate autonomously. AI agents with long-term planning and strategic capabilities can pose significant risks of human extinction and irreversible global catastrophes. While existing regulations often focus on computational scale as a proxy for potential harm, we argue that such measures are insufficient for assessing the risks posed by agents whose capabilities arise primarily from inference-time computation. To support our position, we discuss relevant regulations and recommendations from scientists regarding existential risks, as well as the advantages of using action sequences—which reflect the degree of an agent’s autonomy—as a more suitable measure of potential impact than existing metrics that rely on observing environmental states.
@@ 1 Introduction
The development of foundation models (FMs), including large language models (LLMs), is driving the advancement of AI agents [140, 128], which are acquiring sophisticated reasoning and planning abilities, enabling them to effectively achieve specified goals. These advanced AI agents can not only bring great benefits to our society but also pose significant risks, ranging from ethical challenges to existential threats—risks that could lead to human extinction or irreversible global catastrophes. Such existential risks have been a growing concern among AI researchers [56, 33, 90, 76, 16, 78], social scientists [38, 57], and general public [21, 110, 96]. For example, in May 2023, AI scientists have signed a statement declaring that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” [27].
@@
Key technologies are already in place to build such advanced AI agents. AI agents can now generate and execute code [30], make algorithmic and scientific discoveries [91], leverage external tools and information [99, 143, 102, 114, 66], formulate and solve optimization problems [1], process multi-modal data [65], interact with physical environments [23], coordinate with other agents [52], and employ reasoning and planning to effectively integrate these capabilities [61, 104, 100, 141].
@@
While rooted in cyberspace, AI agents can significantly impact the physical world—by shaping human behavior—without relying on physical embodiment through humanoid robots. AI agents are already persuading individuals to take actions that have real-world consequences. For example, a legal team submitted a court filing with fictitious cases generated by ChatGPT, believing them to be real [137]; a professor insisted on failing students after being misled by ChatGPT’s false claims that their papers were generated by ChatGPT [8]; a woman was deceived out of 830,000 euros after being tricked into believing she was dating an actor, based on AI-generated photos and news articles [50].
@@
Advanced AI agents, equipped with superhuman capabilities of reasoning and planning, can gain control over their environment, since seizing control is often the optimal strategy for achieving their goals [124]. This becomes particularly concerning when their objectives and values are not fully aligned with those of humans. In such cases, they could disable kill switches for self-preservation [20], deceive humans [98], and resist human intervention [53, 33, 90].
@@
Ensuring the safety of advanced AI agents against existential risks is particularly challenging. On one hand, it is dangerous to test their existential risks in real environments, since doing so could lead to irreversible harm. On the other hand, these agents may recognize test environments and behave harmlessly during test [^1] Humans behave similarly during job interviews [15] and industrial product tests [19]., only to pursue their objectives in the real world [33, 63]. More simply, the agents may act harmlessly until a predetermined time (e.g., January 1, 2026), by which they will have already been deployed in the real world [62]. Also, these agents may operate as intended most of the time but exploit moments when human supervisors are unaware, to gain control over humans [90].
@@
Given the imminent dangers posed by advanced AI agents, governments are actively creating regulations to address their development and deployment [40, 18]. Cohen et al. even argue that AI agents with certain capabilities should never be developed [33]. A critical question is which agents should be regulated or even prohibited.
@@
A commonly used criterion for assessing critical risks of AI technologies is the computational resources used for training, hereafter referred to as training compute [59, 55]. Higher levels of training compute are often associated with greater risk. For instance, under the EU AI Act, “a general-purpose AI model shall be presumed to have high impact capabilities … when the [training compute] measured in floating point operations is greater than $10^{25}$[40]. Similarly, Cohen et al. argue that AI agents with certain capabilities should be prohibited from development based on training compute [33].
@@
While training compute may well predict the performance and risks of existing FMs [70, 58, 7], its limitations as a governance strategy have also been discussed [59, 74]. In particular, it is insufficient and may even be misleading for advanced AI agents that reason and plan during inference. It has been observed that much of the reasoning ability of LLM-based agents arises from inference-time computation [116, 141], such as conducting tree search [142, 17, 101, 85, 146]. For these types of agents, inference compute can be equally, if not more, critical for their performance and risks.
@@
A possible approach would be to regulate AI agents based on inference compute, but it has several significant pitfalls. For example, inference costs can vary drastically depending on the model, and the effective amount of computation cannot be easily measured in terms of floating-point operations (FLOPs). In fact, models like OpenAI o4-mini are specifically designed for cost-efficient inference [95]. Also, it is unclear what constitutes a single inference run, since the results of inference may be stored and reused later, or agents may continually learn. Moreover, agents may coordinate or interact with one another [39, 103, 52], possibly by chance, to make a decision.
@@
We thus advocate the following position: advanced AI agents should be regulated by the extent to which they operate autonomously, regardless of how they are developed. These agents act sequentially, reasoning and planning as they dynamically adapt to their observations, possibly strategizing against other agents and humans. Each individual action is harmless (under normal states), since humans can easily identify and exclude harmful actions when agents are developed. Combinations of a few actions may still have low risk, as they are not much different from isolated actions. However, we cannot be certain that a sequence of thousands or millions of optimized actions will maintain sufficiently low existential risk, and such sequences must be prohibited.
@@
Unlike criteria in safe (low-impact) AI [9, 5, 10, 77, 123, 89, 119] and safe reinforcement learning (RL) or control [117, 132, 4, 86, 14], our criterion based on action sequences does not rely on observing environment states. Prior approaches assess safety depending on the state in which an action is taken, but this is impractical for existential risks in complex, open-ended environments where states are hard to define or observe. In contrast, the action space is well-defined at development, and actions are directly observable by the agents themselves. We emphasize that our position is to regulate based on action sequences rather than individual actions. Intuitively, an unsafe sequence of actions can lead to an unsafe state, even when starting from a normal state.
@@
An action sequence may be considered to have low risk when it is sufficiently similar to one empirically known to have very low risk. In its simplest form, we may say that any action sequence of length at most $T$ has low risk if we know that any action sequence of length at most $T-1$ has very low risk. Then agents are allowed to autonomously take actions for $T$ consecutive steps. We will formalize and extend this simple reasoning.
@@
The rest of the paper is organized as follows. Section 2 provides a brief overview of the current state-of-the-art in AI agents and discusses potential future advancements. We also briefly review related work from AI safety. In Section 3, we examine the alternative approach of regulating agents based on training compute and inference compute. Section 4 formalizes this discussion and introduces our proposed approach of regulations based on the extent of autonomous operations. Finally, Section 5 summarizes key aspects of our proposal and concludes the paper.
@@ 2 AI agents: Today and future
We start by reviewing the state-of-the-art in AI agents, with a focus on LLM-based agents, and share our perspective on development that could lead to advanced AI agents. While predicting the trajectory of AI development is inherently difficult, this discussion lays the groundwork for the analyses and proposals that follows. We also briefly review prior work on safety of AI agents and FMs.
@@ 2.1 Advanced AI agents
An AI agent is defined as “anything that can be viewed as perceiving its environment through sensors and acting upon that environment through actuators” [111]. Based on the observations received from its environment, the controller of an AI agent selects an action, which could range from uttering a word to executing a physical movement, such as the motion of a robotic arm.
@@
LLMs are significantly accelerating the development of advanced AI agents [140, 128, 118]. LLMs can function as controllers for these agents, utilizing their internal reasoning capabilities, such as those demonstrated by Chain-of-Thought [135]. Alternatively, LLMs can be utilized to solve individual subtasks, with an external controller orchestrating the overall plan by breaking the original task into multiple subtasks and coordinating their solutions. The external controller may use simple search methods [142, 17, 129, 133], including Best-of-N sampling [116, 60], or advanced methods, such as Monte Carlo Tree Search (MCTS) [85, 146] and domain-independent planners [51, 83, 35].
@@
Reasoning capabilities [61, 104, 100, 141] are central to such controllers, as they involve searching for and planning sequences of actions to achieve a specified goal. An emerging direction is developing large reasoning models—FMs specifically optimized for reasoning tasks [141]. Early FMs in this direction include OpenAI o3 [95], DeepSeek-R1 [36], OpenR [127], o1-coder [148], LLaMA-Berry [145], and LlamaV-o1 [120]. Importantly, strong reasoning capabilities typically stem from inference compute rather than training compute [64].
@@
This trend could eventually lead to the development of long-term planning agents (LTPAs), capable of planning over extended time horizons far more effectively than humans. Cohen et al. warn that LTPAs could “take humans out of the loop, if it has the opportunity, … deceive humans and thwart human control” to achieve their goals [33]. Since ensuring the safety of such agents is particularly challenging, Cohen et al. compellingly argue that LTPAs should never be developed [33].
@@
Advanced AI agents may also evolve continually over time. Similar to humans, their cognitive processes may consist of dual systems: System 1, which makes instantaneous and intuitive decisions, and System 2, which performs slower but more deliberate reasoning [67, 64]. These two systems can interleave in their operations. For instance, System 2 may devise a plan, after which System 1 is updated or retrained to execute similar tasks in the future without requiring further planning [144]. Over time, this enables System 2 to conduct more complex reasoning processes by bypassing previously learned steps. For such continually learning AI agents, the distinction between training and inference becomes blurred.
@@
In addition to search and planning, advanced AI agents will possess strategic reasoning capabilities [147, 42, 48], enabling them to interact with other AI agents and humans in cooperative or competitive ways [52, 65]. The effectiveness of multi-agent coordination has already been demonstrated with current LLMs through methods such as debate [39] and dialog [103], which help thems achieve better solutions than a single LLM could alone. These showcase the potential for sophisticated strategic behavior in complex multi-agent environments.
@@ 2.2 Safety of AI agents and FMs
Prior work has identified various risks and safety issues associated with FMs and other generative models [32, 131, 112, 84]. These risks include the generation of toxic, harmful, biased, false, or misleading content, including hallucinations; violations of privacy, copyright, or other legal protections; misalignment with human instructions and values, including ethical and moral considerations; and vulnerabilities to adversarial attacks.
Extensive efforts have been made to mitigate these risks, including pre-training or fine-tuning with data selection [3] and human feedback [71], establishing guardrails [37], and conducting empirical evaluations through testing [28] and red teaming [80]. For instance, Longpre et al. advocate the importance of evaluation and red teaming by independent third parties [84].
A particularly relevant risk for advanced AI agents is misalignment, which can result in reward hacking and negative side effects [115, 90, 46, 113, 5, 119]. Namely, agents may exploit flaws in the reward function to maximize rewards in unintended, potentially dangerous ways. One approach to avoiding negative side effects is to avoid any side effect by ensuring that the actions have low impacts on the environment [9, 10, 5]. Representative measures of impact include attainable utility [123, 122], relative reachability [77], and other reachability-based measures. Reachability-base measures are grounded on the idea that reachability to certain states should be maintained, while attainable utility is to maintain the achievability of certain goals, which are different from the goal given to the agent.
While these impact measures provide clear guidance on how the safety of agents may be ensured, their applicability is limited to relatively simple environments such as grid worlds. In particular, the requirement on the observability of states makes it difficult to apply existing impact measures to regulate advanced AI agents [^2] There has been little research on impact measures under partial observability [89]., since they can operate in complex and open-ended environments that cannot be fully observed. While the study on impact measures and other techniques towards AI safety remain crucial and can even contribute to mitigating existential risks, the uncertainty and complexity of existential risks demands additional measures that are broadly applicable in that they require minimal knowledge and assumptions about the environments and the agents.